OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 − 25 = 0 and x2 = 2x + 15 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points)

OpenStudy (anonymous):

I did Part 1, and 2 I just need help with Part 3. Part 1.) x^2 - 25 + 25 = 0 + 25 x^2 = 2x + 15 x^2 = 25 x^2 - (2x + 15) = 2x + 15 - (2x + 15) x = +sqrt(25) x^2 - 2x - 15 = 0 Answer: x = 5, -5 (x + 3)(x - 5) = 0 x + 3 = 0 or x - 5 = 0 Answer: x = -3 or x = 5 Part 2.) Both of the solution represents y values to equal zero and the only two points in the x-axis in which the functions crosses.

OpenStudy (anonymous):

@Destinycholie

OpenStudy (anonymous):

I said part 3.... @Destinycholie

OpenStudy (anonymous):

@Jlynn89 @haleyelizabeth2017 @sammixboo can you help me pls

OpenStudy (anonymous):

Part 3: How are the graphs alike? How are they different? (2 points)

sammixboo (sammixboo):

Is there graphs given? If so can I see them :D

OpenStudy (anonymous):

No sorry I think I have to make my own but all I need is how do they look alike and how their different

OpenStudy (anonymous):

I can make a graph really quick

sammixboo (sammixboo):

OKay thanks! Here is some things you can ask yourself Are the Chart Names alike or different? Is all the x and y label on the graph the same or different? Do they both go to the same limit? Etc...